40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 252     Entry time: Tue Jan 22 02:33:45 2008
Author: rob 
Type: Update 
Category: LSC 
Subject: DRMI work 

0) The ETMY oplev needs work/centering

1) recentered DRMI oplevs

2) Did some light DRMI locking. Looked at the loops and the DD signals. The PODD signals look flaky; the beam may not be on the diode. MICH and PRC handoffs to DD signals were spotty, but not a total disaster. Changed the PD9 phase by 115 degs. Work continues on the DD_handoff subscript.

3) John says "There are ants everywhere."

4) Andrey is now versed in the arts of decimation.
ELOG V3.1.3-